OpenStudy (anonymous):

Perform the indicated operations. If the matrix does not exist, write impossible. [8 2 6] + [11 -7 1] Please help & explain ((:!

OpenStudy (anonymous):

Matrix addition is done entry by entry. In general:\[\begin{bmatrix} a & b & c \end{bmatrix} + \begin{bmatrix} d & e & f \end{bmatrix} = \begin{bmatrix} a + d & b + e & c + f \end{bmatrix}\]

OpenStudy (anonymous):

Can you explain to me & answer. I am confused.

OpenStudy (anonymous):

What are a, b, c, d, e and f in your problem looking at my comment above?

OpenStudy (anonymous):

[-8 2 6] + [11 -7 1] = [-8 + 11 2 + -7 6+1]

OpenStudy (anonymous):

That is correct. So... -8 + 11 = ? 2 + -7 = ? 6+1 = ?

OpenStudy (anonymous):

-8 + 11 = 3 2 + -7 = -5 6 + 1 = 7